OpenStudy (anonymous):

help...Domingo works 40 hours per week as a tire store manager. If he made $29,640 last year, how much was he paid per hour?

jimthompson5910 (jim_thompson5910):

Let x = hourly wage If he worked 40 hours in one week, then he earned 40x dollars in one week Assuming he works 52 weeks a year, and he earns $29,640 in a year, then this means 52*40x = 29640 2080x = 29640 x = 29640/2080 x = ???

OpenStudy (anonymous):

$14.25!

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

x = 29640/2080 x = its 14.25 once u find x